Sheridan provides environmentally friendly lawn services for homeowners. Its operating costs are as follows.

Depreciation $1,400 per month
Advertising $250 per month
Insurance $2,475 per month
Weed and feed materials $17 per lawn
Direct labor $8 per lawn
Fuel $2 per lawn


Sheridan charges $60 per treatment for the average single-family lawn.

1. Determine the company’s break-even point in number of lawns serviced per month.

2. Determine the company’s break-even point in dollars.
Answer #1

Answer

Fixed cost = 1,400 + 250 + 2,475 = $4,125

Total variable cost per unit = 17 + 8 + 2 = $27

Selling price = $60

Contribution margin = selling price - variable cost

Contribution margin = 60 - 27 = $33

Contribution margin ratio = contribution margin / selling price

Contribution margin ratio = 33 / 60 = .55

1 Break even point in unit = fixed cost / contribution margin

Break even point in unit = 4,125 / 33 = 125 units

2 Break even point in dollar = fixed cost / contribution margin ratio

Break even point in dollar = 4,125 / .55 = $7,500

The company break even point in unit is 125 or in dollar is $7,500.

Fixed cost includes depreciation , advertising and insurance and variable cost includes weed and feed materials, direct labour and fuel.

If YOU are like the answer please upvote.